All answers must be entered as formulas (using cell names)AB Makers Corp. had additions to retained earnings for the year just ended of $395,000. The firm paid out $195,000 in cash dividends, and it has ending total equity of $5.3 million. If the 3company currently has 170,000 shares of common stock outstanding, what are earnings per share? Dividends per share? Book value per share? If the stock currently sells for $64 per share, what is the market-to-book ratio? The price-earnings ratio? If the company had sales of $5.15 million, what is the price-sales ratio? 4 6 Addition to retained earnings 7Cash dividends 395,000 $195,000 $5,300,000 170,000 64 $5,150,000 8 Total equity 9 Common shares outstanding 10 Share price 1Sales 12 14 15 16 17 18 Complete the following analysis. Do not hard code values in your calculations. Net income Earnings per shareAB 13 14 Complete the following analysis. Do not hard code values in your calculations. 15 16Net income 17 18Earnings per share 19 20Dividends per share 21 22Book value per share 23 24Market-to-book ratio 25 26Price-earnings ratio 27 28Sales per share 29 30 Price-sales ratio 31 32 34
Answer #1

1) Net Income = Dividend distributed + Addition made to reserves

=195000+395000

=$590,000

2) EPS = Net Income/No of shares

=590,000/170,000

=$3.47

3) Dividend per share = 195000/170,000

=$1.147

4)Book value per share = Total Equity/No of shares

=5300000/170000

=$31.176

5) Market to book ratio = Market price per share/Book value per share

=64/31.176

=2.05 times

6) Price earning ratio = Market price per share/Earning per share

=64/3.47

=18.44

7) Sales per share = Total sales/No of shares

=5150000/170000

=30.294

8) Price- sales ratio = (No of shares * Market price per share)/ Sales

=(170000*64)/5150000

=2.11